The following contingency table represents the voters’ opinion concerning a new tax



reform bill and their level of income.



Levels of Income



Opinion Low Medium High



For 213 203 182



Against 138 110 154



Apply chi-square (χ2



) test to examine whether the opinion of people on new tax bill is



independent from their income level. [Table value of chi- square (χ2



) at 5% level of



significance and for 2 degrees of freedom is 5.99]



1
Expert's answer
2022-01-09T10:37:59-0500

"H_0:" the opinion of people on new tax bill is independent from their income level

"H_a:" ​the opinion of people on new tax bill is dependent from their income level


"\Chi^2=\sum \frac{(O_i-E_i)^2}{E_i}=7.878"

where

Oi is observation value,

Ei is expected value

Ei = row total * column total/ total


"df=(Rows-1)*(Columns-1) = (2-1)*(3-1) = 2"

critical value:

"\Chi^2_{crit}=5.99"


Since "\Chi^2>\Chi^2_{crit}" we reject the null hypothesis. ​The opinion of people on new tax bill is dependent from their income level.